A balloon is filled with helium and let go. It rises into the air accelerating at a rate of 0.151 m s−2.

If the density of air is 1.22 kg m−3 then what is the average density of the balloon? (to 3 s.f and in kg m−3)

(Note: use g = 9.81 m s−2 and ignore any "drag" forces on the balloon such as air resistance.)
1
Expert's answer
2020-04-23T12:46:24-0400
"ma=\u03c1 _\na\n\u200b\t\n gV\u2212mg"

"\u03c1 _\nb\n\u200b\t\n Va=\u03c1 \n_a\n\u200b\t\n gV\u2212\u03c1 \n_b\n\u200b\t\n gV"

"\u03c1 \n_b\n\u200b\t\n (a+g)=\u03c1 _\na\n\u200b\t\n g"

"\u03c1 _\nb\n\u200b\t\n (0.151+9.81)=1.22(9.81)"

"\u03c1 _\nb\n\u200b\t\n =1.20\\frac{kg}{m^3}"